# NeuroLifting: Neural Inference on Markov Random Fields at Scale

### v2

@@ 6 Conclusion
In this paper, we introduced NeuroLifting and its application to solving MAP problems for MRFs. Our experiments showed that NeuroLifting effectively handles MRFs of varying orders and energy functions, achieving performance on par with established benchmarks, as verified on the UAI 2022 inference competition dataset. Notably, NeuroLifting excels with large and dense MRFs, outperforming traditional methods and competing approaches on both synthetic large instances and real-world PCI instances.This method, which utilizes Neural Networks for lifting, has proven successful and could potentially be extended to other optimization problems with similar modeling frameworks.
@@ Appendix A Padding procedure
Appendix B Limitations
Our proposed GNN-based approach, while effective for complex MRF problems, presents several limitations worth acknowledging. The method’s computational overhead makes it less efficient for small instances where traditional algorithms may perform adequately without the preprocessing and inference costs of neural networks. Additionally, memory requirements for maintaining graph structures during message passing can become prohibitive for extremely large MRFs. Future work should focus on fully fully leveraging the potential of our method and balancing the trade-off between powerful representations and computational efficiency.
Appendix C Full results on Pairwise Synthetic instances
@@
Appendix D Results of UAI Inference Competition 2022 dataset
@@
Table 5: Results on the UAI inference competition 2022. Numbers correspond to the obtained energy values. Best in bold.“opt" denotes it is the optimal solution.
@@
Maxsat_mod4block_2vars_10gates_u2_autoenc 479/123509 -$1$186103.111 146166.620 $\pm$ 41250.035 -146166.797 $\pm$ 41249.859
@@
Appendix E More analysis
@@ Appendix F GNN formulations
Appendix G Results about the analysis experiments
@@
Figure 4: The landscape of instance Segmentation_19. From top to the bottom, each column correspond to network layer $\{1,2,5,8\}$. The first row is the landscape range from $[-10,+10]$ for both $\delta$ and $\eta$ direction. The second row is the landscape range from $[-1,+1]$ for both $\delta$ and $\eta$ direction.
@@
Table 9: Time comparison between Toulbar2 and NeuroLifting on PCI instances."-" if the solving process is already stopped.
@@
Appendix H Visualization setup
The core idea of the visualization technique proposed by visualloss [59] involves applying perturbations to the trained network parameters $\theta^{*}$ along two directional vectors, $\delta$ and $\eta$: $f(\alpha,\beta)=L(\theta^{*}+\alpha\delta+\beta\eta)$. By doing so, we can generate a 3-D representation of the landscape corresponding to the perturbed parameter space.
We sampled 250000 points in the $\alpha-\beta$ plane, where both $\alpha$ and $\beta$ range from -10 to 10, to obtain an overview of the loss function landscape. Subsequently, we focused on the region around the parameter $\theta^{*}$ by sampling an additional 10,000 points in a narrower range, with $\alpha$ and $\beta$ both from $-1$ to $1$.
Appendix I Read UAI format files
An example data file in UAI format is provided in Box. I. This Markov Random Field consists of 3 variables, each with 2 possible states. Detailed information can be found in the box, where we illustrate the meanings of different sections of the file. Notably, in the potential section, the distributions are not normalized. During the BP procedure, these distributions will be normalized to prevent numerical issues. However, in the energy transformation phase, we will utilize these values directly.
@@ Appendix I Read UAI format files
1 2 2 0 1     //2 means this clique is an edge, the edge is (0, 1).
@@
0.1 10 2 0.5 0.5 4 0.1 1.0 1.0 0.1//The potential of the state combinations for variables 0 and 1 is given in the order of (0,0), (0,1), (1,0) and (1,1).
@@
Appendix J PCI problem formulation
@@
$\min\limits_{z,L}\quad$ $\sum\limits_{(i,j)\in\mathcal{E}}a_{ij}L_{ij}$ (9)
$\mathrm{s.t.}\quad$ $z_{np}\in\{0,1\},\quad\forall n\in N,p\in P$ (9a)
$\sum\limits_{p\in P}z_{np}=1,\quad\forall n\in N.$ (9b)

### v3

@@ 6 Conclusion
In this paper, we introduced NeuroLifting for solving MAP problems in MRFs. Our experiments demonstrate that NeuroLifting effectively handles MRFs of varying orders and energy functions, achieving competitive performance on the UAI 2022 dataset. Notably, it excels with large and dense MRFs, outperforming traditional and competing methods on both synthetic and real-world PCI instances. This neural lifting approach could potentially extend to other optimization problems with similar frameworks.
Impact Statement
This work presents a novel unsupervised framework designed for inference tasks on arbitrary MRFs. Our proposed method achieves comparable performance to exact solvers on tractable instances while demonstrating superior scalability on large-scale problems that exceed the computational capabilities of existing approaches, including both traditional belief propagation methods and contemporary deep learning techniques. The framework’s versatility and scalability make it particularly valuable for practical applications of MRF inference across various industrial domains.
@@ Appendix A Padding procedure
Appendix B Proofs of Section. 4.5
B.1 Proof of Proposition 4.3
[…]
Proof.
As each $\Delta_{i}$ is a convex polytope, the Cartesian product $\mathcal{P}$ is also a convex polytope. For a node $k$, we can reformulate the relaxed energy as:
$\hat{E}(h_{k},H_{-k})=\underbrace{\sum_{s\in S_{k}}h_{k,s}\theta_{k}(s)}_{\text{unary term}}+\underbrace{\sum_{C:k\in C}\sum_{y_{C}}\left(h_{k,y_{k}}\prod_{j\in C,j\neq k}h_{j,y_{j}}\right)\theta_{C}(y_{C})}_{\text{clique term}}+\text{const}$ (11)
We can rewrite this into a linear format w.r.t. $h_{k}$:
$\hat{E}(h_{k})=\sum_{s\in S_{k}}h_{k,s}\cdot\underbrace{\left(\theta_{k}(s)+\sum_{C:k\in C}\sum_{y_{C\backslash k}}\left(\prod_{j\in C,j\neq k}h_{j,y_{j}}\right)\theta_{C}(s,y_{C\backslash k})\right)}_{\text{coefficient}}+\text{const}$ (12)
This can be further simplified:
$\hat{E}(h_{k})=\langle h_{k},\mathbf{C}_{k}\rangle+\text{const}$ (13)
where $\mathbf{C}_{k}$ is a coefficient vector depending only on connecting nodes $H_{-k}$. Thus, for any variable $h_{k}$, $\hat{E}(H)$ is a linear function (multilinear).
Consider $\min_{h_{k}\in\Delta_{k}}\hat{E}(h_{k})$, which is a linear programming problem on the simplex $\Delta_{k}$. According to the fundamental theorem of linear programming, the minimum must lie on a vertex. Thus, for any given $H_{-k}$, there exists a one-hot vector $h_{k}^{*}$, such that $\hat{E}(h_{k}^{*},H_{-k})\leq\hat{E}(h_{k},H_{-k})$. Consequently, the global minimum of $\hat{E}(H)$ must be obtained at a vertex $H^{*}$ of $\mathcal{P}$. Since the vertices of $\mathcal{P}$ are isomorphic to $\mathcal{X}$, we have:
$\min_{H\in\mathcal{P}}\hat{E}(H)=\min_{H\in\mathcal{V}(\mathcal{P})}\hat{E}(H)=\min_{x\in\mathcal{X}}E(x)$ (14)
B.2 Proof of Theorem 4.5
[…]
Proof.
We construct a continuous, energy-bounded path through three phases:
1. Retraction Phase ($t\in[0,1/3]$): The parameters are scaled linearly from $W_{a}$ to a low-norm regime $\epsilon W_{a}$. According to Assumption 4.4, the distribution $H$ moves from $H^{(a)}$ toward the uniform distribution $U$. Because $\hat{E}$ is multilinear with respect to each node’s probabilistic assignment, the energy along this segment is monotonically bounded by $\max(E(H^{(a)}),\hat{E}(U))$.
2. Rotation Phase ($t\in[1/3,2/3]$): The parameter direction is rotated from $\hat{W}_{a}$ to $\hat{W}_{b}$ while keeping $W||\approx\epsilon$. In the original discrete space $\mathcal{X}$, a barrier is a codimension-1 hypersurface that separates minima. However, by Assumption 4.5, the over-parameterized space $\mathcal{W}$ provides “extra room”. By transversality arguments, there exists a path within a tubular neighborhood of the fiber $\Phi^{-1}(U)$ that bypasses these ridges. Thus, the maximum energy in this phase is bounded by $\hat{E}(U)+\delta$.
3. Expansion Phase ($t\in[2/3,1]$): The norm is increased from $\epsilon$ toward the target magnitude $W_{b}$. The Softmax function sharpens the distribution, moving $H$ from $U$ to the basin of $H^{(b)}$. The energy remains bounded by $\max(\hat{E}(U),E(H^{(b)}))$.
The piecewise continuity of $W(t)$ and the existence of a high-dimensional bypass demonstrate that GNN parameterization constructs a barrier-free trajectory, which further eases the optimization dynamics of gradient descent. ∎
B.3 Proof of Theorem 4.6
[…]
Proof.
If $\mathcal{L}$ is $L_{total}$-Lipschitz, there for any $W$ and $W^{\prime}$, we have:
$\mathcal{L}(W^{\prime})\leq\mathcal{L}(W)+\langle\nabla\mathcal{L}(W),W^{\prime}-W\rangle+\frac{L_{total}}{2}\|W^{\prime}-W\|^{2}$ (15)
We let $W_{t}=W$ be the current parameters and $W_{t+1}=W^{\prime}$ be the updated parameters after one round.
Substituting $W_{t+1}=W_{t}-\eta\nabla\mathcal{L}(W_{t})$ into the inequality above. The displacement is:
$W_{t+1}-W_{t}=-\eta\nabla\mathcal{L}(W_{t})$
The linear term becomes:
$\langle\nabla\mathcal{L}(W_{t}),-\eta\nabla\mathcal{L}(W_{t})\rangle=-\eta\|\nabla\mathcal{L}(W_{t})\|^{2}$
The quadratic term is:
$\frac{L_{total}}{2}\|-\eta\nabla\mathcal{L}(W_{t})\|^{2}=\frac{L_{total}\eta^{2}}{2}\|\nabla\mathcal{L}(W_{t})\|^{2}$
Thus:
$\mathcal{L}(W_{t+1})\leq\mathcal{L}(W_{t})-\eta\|\nabla\mathcal{L}(W_{t})\|^{2}+\frac{L_{total}\eta^{2}}{2}\|\nabla\mathcal{L}(W_{t})\|^{2}=\mathcal{L}(W_{t})-\eta\left(1-\frac{\eta L_{total}}{2}\right)\|\nabla\mathcal{L}(W_{t})\|^{2}$
To guarantee the decrease of $\mathcal{L}$, the second term of the righthand side of the inequality is non-negative:
$\eta\left(1-\frac{\eta L_{total}}{2}\right)>0$
As $\eta>0$, we only need:
$1-\frac{\eta L_{total}}{2}>0\Rightarrow\eta<\frac{2}{L_{total}}$
This will lead to monotonic non-decrease of the energy.
Summing the inequality from $t=0$ to $T$:
$\sum_{t=0}^{T}(\mathcal{L}(W_{t})-\mathcal{L}(W_{t+1}))\geq\sum_{t=0}^{T}\delta\|\nabla\mathcal{L}(W_{t})\|^{2}$
As the energy has a lower bound $\mathcal{L}^{*}$ (since MRF is defined on a finite discrete space), we have:
$\mathcal{L}(W_{0})-\mathcal{L}^{*}\geq\mathcal{L}(W_{0})-\mathcal{L}(W_{T+1})\geq\sum_{t=0}^{T}\delta\|\nabla\mathcal{L}(W_{t})\|^{2}$
Indicating the convergence of the series. Thus we have:
$\lim_{t\rightarrow\infty}\|\nabla\mathcal{L}(W_{t})\|=0$
This means that the optimization of $\hat{E}(H)$ will converge to a stationary point. ∎
Appendix C Complexity Analysis
The primary computations in this model arise from both the loss calculation and the operations within the GNN. For the loss function, let $c_{max}$ denote the maximum clique size. The time complexity for the loss calculation is given by $O(|\mathcal{V}||\mathcal{X}|+c_{max}|\mathcal{C}||\mathcal{X}|)$. F or the GNN component, let $\mathcal{N}_{v}$ denote the average number of neighbors per node in the graph. The time complexity for neighbor aggregation in each layer is $O(\mathcal{N}_{v}|\mathcal{V}|)$ , and merging the results for all nodes requires $O(|\mathcal{V}|d)$ where $d$ is the feature dimension. Thus, for a $K$-layer GraphSAGE model with the custom loss function, the overall time complexity can be expressed as $O(|\mathcal{X}|(|\mathcal{V}|+c_{max}|\mathcal{C}|)+K|\mathcal{V}|(\mathcal{N}_{v}+d))$. This analysis highlights the efficiency of the framework in managing large-scale graphs by leveraging neighborhood sampling and aggregation techniques. The derived complexity indicates that the model scales linearly with respect to the number of nodes, the number of layers, and the dimensionality of the feature vectors, making it well-suited for large-scale instances.
Appendix D Full results on Pairwise Synthetic instances
@@
Appendix E Results of UAI Inference Competition 2022 dataset
@@
Table 5: Results on the UAI inference competition 2022. Numbers correspond to the obtained energy values. Best in bold.“opt” denotes it is the optimal solution.
@@
Maxsat_mod4block_2vars_10gates_u2_autoenc 479/123509 -186103.111 -186103.111 146166.620 $\pm$ 41250.035 -146166.797 $\pm$ 41249.859
@@
Appendix F More analysis
@@ Appendix G GNN formulations
Appendix H Results about the analysis experiments
@@
Figure 4: The landscape of instance Segmentation_19. From top to the bottom, each column correspond to network layer $\{1,2,5,8\}$. The first row is the landscape range from [10,+10][-10,+10] for both $\delta$ and $\eta$ direction. The second row is the landscape range from [1,+1][-1,+1] for both $\delta$ and $\eta$ direction.
@@
Table 9: Time comparison between Toulbar2 and NeuroLifting on PCI instances. “-” if the solving process is already stopped.
@@
Appendix I Visualization setup
The core idea of the visualization technique proposed by Li et al. (2018) involves applying perturbations to the trained network parameters $\theta^{*}$ along two directional vectors, $\delta$ and $\eta$: $f(\alpha,\beta)=L(\theta^{*}+\alpha\delta+\beta\eta)$. By doing so, we can generate a 3-D representation of the landscape corresponding to the perturbed parameter space.
We sampled 250000 points in the $\alpha-\beta$ plane, where both $\alpha$ and $\beta$ range from -10 to 10, to obtain an overview of the loss function landscape. Subsequently, we focused on the region around the parameter $\theta^{*}$ by sampling an additional 10,000 points in a narrower range, with $\alpha$ and $\beta$ both from 1-1 to $1$.
Appendix J Read UAI format files
An example data file in UAI format is provided in Box. J. This Markov Random Field consists of 3 variables, each with 2 possible states. Detailed information can be found in the box, where we illustrate the meanings of different sections of the file. Notably, in the potential section, the distributions are not normalized. During the BP procedure, these distributions will be normalized to prevent numerical issues. However, in the energy transformation phase, we will utilize these values directly.
@@ Appendix I Visualization setup
1 2
2 0 1    //2 means this clique is an edge, the edge is (0, 1).
@@
0.1 10
2
0.5 0.5
4
0.1 1.0 1.0 0.1//The potential of the state combinations for variables 0 and 1 is given in the order of (0,0), (0,1), (1,0) and (1,1).
@@
Appendix K PCI problem formulation
@@
$\min\limits_{z,L}\qquad$ $\sum\limits_{(i,j)\in\mathcal{E}}a_{ij}L_{ij}$ (16)
$\mathrm{s.t.}\qquad$ $z_{np}\in\{0,1\},\qquad\forall n\in N,p\in P$ (16a)
$\sum\limits_{p\in P}z_{np}=1,\qquad\forall n\in N.$ (16b)